# Massless Quasiparticles in Bogoliubov-de Gennes Systems

### v1

@@ Appendix B Proof of $[D(\mathcal{P})K,D(g)K_{s(g)}]=0$
Now we verify the equation (33). Suppose $g$ acts on $C_{\uparrow}^{\dagger}$ as
@@ Appendix C Generalized band representation for SCs/SFs/QSLs
When the pairing terms are switched off, the SCs/SFs become metals or insulators where a number of localized orbitals in the unit cell in real space determine the energy band structure in the Brillouin zone. The energy bands as entire entities form a special representation of the system’s symmetry group $G$ called the band representation Zak [1982], Bradlyn et al. [2017b], Cano and Bradlyn [2021]. Such a band rep is called elementary if it cannot be decomposed as a direct sum of smaller band representations. Once the elementary band reps are known, all the band reps can be obtained by the direct sum of elementary ones.
The band representations can be generalized to BdG systems by introducing ‘particle-hole’ degrees of freedom. We start from the full ‘symmetry’ group $F_{f}=G_{f}+\mathcal{P}G_{f}$. The construction of an elementary band representation needs two ingredients. The first is a ‘maximal’ Wyckoff position $q_{1}$ with a site-symmetry group $G_{q_{1}}$ (composed of operations in $G_{f}$ which stabilize the site $q_{1}$). The second is a set of creation and annihilation operators of localized orbitals $\phi_{1}(q-q_{1}-R_{m}),\phi_{2}(q-q_{1}-R_{m}),...,\phi_{r}(q-q_{1}-R_{m})$ centered at $q_{1}$ in the $m$-th unit cell, namely $[c_{\sigma,q_{1}+R_{m}}^{\dagger},c_{\sigma,q_{1}+R_{m}}]$ with $\sigma=1,2,...,r$ and
@@
Moreover, if one defines the representation matrix for the little co-group element $\alpha\in G_{f}(k)$ with
$D^{(l)k}_{(\Sigma,q_{i}),(\Sigma^{\prime},q_{j})}(\alpha)$ $=$ $D^{(l)k}_{(\Sigma,q_{i}),(\Sigma^{\prime},q_{j})}(\alpha)e^{\mathrm{i}k\cdot t_{\alpha}}$
@@
then $D^{(l)k}_{(\Sigma,q_{i}),(\Sigma^{\prime},q_{j})}(G_{f}(k))$ form a projective rep of the little co-group $G_{f}(k)$ with the factor system $\omega_{2}(\alpha,\beta)=e^{-iK_{\alpha}\cdot t_{\beta}}$ where $K_{\alpha}=p_{\alpha}^{-1}\cdot k-k$ and $\alpha,\beta\in G_{f}(k)$.
@@ Appendix D The restriction from irreps of $F_{f}(k)$ to its normal subgroup $G_{f}(k)$
The little co-group at $k$ has the following structure $F_{f}(k)=G_{f}(k)+\tilde{\mathcal{P}}G_{f}(k)$, with
$F_{f}(k)/G_{f}(k)\cong Z_{2},$
where the group $G_{f}(k)$ can be either unitary or anti-unitary, and the same as the effective particle-hole operator $\tilde{\mathcal{P}}$. Denoting $D(F_{f}(k))$as an irreducible representation of $F_{f}(k)$, and denoting $\mathcal{L}$ as its supporting Hilbert space, then we have the following theorem.
@@
The restricted representation $D(G_{f}(k))$ from irrep $D(F_{f}(k))$ contains at most two irreducible representations of $G_{f}(k)$. There are two possibilities:
  • I,
    $\mathcal{L}$ carries an irrep of $G_{f}(k)$ and the $\mathcal{L}_{0}$ is invariant under the action of $\tilde{\mathcal{P}}$.
  • II,
    $\mathcal{L}$ is reducible for $G_{f}(k)$, namely $\mathcal{L}=\mathcal{L}_{\varepsilon}\oplus\mathcal{L}_{-\varepsilon}$ where each of the two subspaces $\mathcal{L}_{\varepsilon}$ and $\mathcal{L}_{-\varepsilon}$ carries an irrep of $G_{f}(k)$. The coset element $\tilde{\mathcal{P}}$ permutes the two subspaces $\mathcal{L}_{\varepsilon},\mathcal{L}_{-\varepsilon}$.
@@
Since $\mathcal{L}$ carries a representation of $G_{f}(k)$, it must contain an invariant subspace $\mathcal{L}_{\varepsilon}$ of $G_{f}(k)$. Namely, for any $g\in G_{f}(k)$, one has
$g\mathcal{L}_{\varepsilon}=\mathcal{L}_{\varepsilon}.$
@@
Meanwhile, $\mathcal{L}$ carries an irrep of $F_{f}(k)$, so $\tilde{\mathcal{P}}$ transforms $\mathcal{L}_{\varepsilon}$ into another linear subspace $\mathcal{L}_{-\varepsilon}\equiv\tilde{\mathcal{P}}(\mathcal{L}_{\varepsilon})$:
$h_{k}\left(\tilde{\mathcal{P}}\mathcal{L}_{\varepsilon}\right)$ $=$ $\tilde{\mathcal{P}}\left(\tilde{\mathcal{P}}^{-1}h_{k}\tilde{\mathcal{P}}\right)\mathcal{L}_{\varepsilon}$
$=$ $-\tilde{\mathcal{P}}h_{k}\mathcal{L}_{\varepsilon}=-\varepsilon\left(\tilde{\mathcal{P}}\mathcal{L}_{\varepsilon}\right).$
Then it is easily seen that $\mathcal{L}_{-\varepsilon}$ is also invariant under the action of $g\in G_{f}(k)$:
$g\mathcal{L}_{-\varepsilon}=g\tilde{\mathcal{P}}\mathcal{L}_{\varepsilon}$ $=$ $\tilde{\mathcal{P}}\left(\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}}\right)\mathcal{L}_{\varepsilon}$
$=$ $\tilde{\mathcal{P}}g^{\prime}\mathcal{L}_{\varepsilon}=\tilde{\mathcal{P}}\mathcal{L}_{\varepsilon}=\mathcal{L}_{-\varepsilon},$
@@
where we have used the fact that $g^{\prime}=(\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})\in G_{f}(k)$. Therefore, $\mathcal{L}_{-\varepsilon}$ also carries an irrep of $G_{f}(k)$.
@@
Since both $\mathcal{L}_{\varepsilon}$ and $\mathcal{L}_{\varepsilon^{\prime}}$ are irreducible representation spaces of $G_{f}(k)$, there are two possibilities:
(I) $\mathcal{L}_{\varepsilon}=\mathcal{L}_{-\varepsilon}=\mathcal{L}$. In this case, $\varepsilon=-\varepsilon$, meaning that $\mathcal{L}$ forms a set of irreducible zero modes.
(II) $\mathcal{L}_{\varepsilon}\cap\mathcal{L}_{-\varepsilon}=\emptyset$. In this case, $\mathcal{L}=\mathcal{L}_{\varepsilon}\cup\mathcal{L}_{-\varepsilon}$, meaning that $\mathcal{L}$ forms a set of irreducible nonzero modes.
In the following we discuss three different situations of $F_{f}(k)$.
(1) $F_{f}(k)$ is a unitary group. The irrep of $F_{f}(k)$ falls into two classes, extension from the subgroup $G_{f}(k)$ and induction from the subgroup $G_{f}(k)$. The first class corresponds to the case I in the theorem5(irreducible zero modes) and the second class corresponds to the case II (irreducible nonzero modes).
@@
(2) $F_{f}(k)$ is an anti-unitary group, $G_{f}(k)$ is unitary and $\tilde{\mathcal{P}}$ is anti-unitary. The irrep of $F_{f}(k)$ falls into three classes, namely the $\mathbb{R},\mathbb{C},\mathbb{H}$ classes. The $\mathbb{R}$ class corresponds to the case I (irreducible zero modes) in the theorem5; the $\mathbb{C}$ and $\mathbb{H}$ classes correspond to the case II (irreducible nonzero modes, the two classes are distinguished by whether the representations carried by $\mathcal{L}_{\varepsilon}$ is equivalent to that of $\mathcal{L}_{-\varepsilon}$ or not)Shaw and Lever [1974].
@@
(3) $F_{f}(k)$ is anti-unitary, $G_{f}(k)$ is anti-unitary and $\tilde{\mathcal{P}}$ is unitary. Similar to the case in which $F_{f}(k)$ is unitary, the irrep of $F_{f}(k)$ can be either extension or induction from the subgroup $G_{f}(k)$.
@@
The structure of the irrep in case I is relatively simple since both $D(F_{f}(k))$ and $D(G_{f}(k))$ are irreducible. In the following we will discuss the structure of the irrep in case II.
Fixing a set of bases $[\Phi]=[\phi_{1},...,\phi_{d}]$ in $\mathcal{L}_{\varepsilon}$ (assuming the dimensionality of $\mathcal{L}_{\varepsilon}$ is $d$), and denote $M(G_{f}(k))$ as the $d$-dimensional representation of $G_{f}(k)$ on $\mathcal{L}_{\varepsilon}$. In the other subspace $\mathcal{L}_{-\varepsilon}$, we simply choose the bases as $\tilde{\mathcal{P}}[\Phi]=[\tilde{\mathcal{P}}\phi_{1},...,\tilde{\mathcal{P}}\phi_{d}]$. Then in the space $\mathcal{L}$ the particle-hole operator $\tilde{\mathcal{P}}$ is represented as
$D(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}=\left(\begin{matrix}0&\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}})M(\tilde{\mathcal{P}}^{2})\\ I_{d}&0\end{matrix}\right)K_{s(\tilde{\mathcal{P}})},$
where $\tilde{\mathcal{P}}^{2}\in G_{f}(k)$. For any $g\in G_{f}(k)$, we have
$D(g)K_{s(g)}=\begin{pmatrix}M(g)&0\\ 0&M^{\prime}(g)\end{pmatrix}K_{s(g)}.$
with
$M^{\prime}(g)={\omega_{2}(g,\tilde{\mathcal{P}})\over\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})}K_{s(\tilde{\mathcal{P}})}M(\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}.$
The Hamiltonian matrix $h_{k}$ commuting with $D(G_{f}(k)$ and anti-commuting with $D(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}$ is simply given as $h_{k}\equiv\begin{pmatrix}\varepsilon_{k}I_{d}&0\\ 0&-\varepsilon_{k}I_{d}\end{pmatrix}$.
@@ Appendix E Reducible minimal nonzero modes
A set of reducible minimal nonzero modes contains two irreducible irreps of $F_{f}(k)$. To ensure that the two irreps can couple with each other to form nonzero modes, the restricted reps of the normal subgroup $G_{f}(k)$ should be equivalent. Now we start with two irreps $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ satisfying the relation
@@
Defining $M\equiv\left[D(\tilde{\mathcal{P}})D^{\prime-1}(\tilde{\mathcal{P}})\right]$, then we have the following lemma,
@@
The unitary matrix $M=\left[D(\tilde{\mathcal{P}})D^{\prime-1}(\tilde{\mathcal{P}})\right]$ commutes with the restricted rep $D(G_{f}(k))$.
@@
$=$ $MD^{\prime}(\tilde{\mathcal{P}})D^{\prime}(h)K_{s(\tilde{\mathcal{P}}h)}$
$=$ $\omega_{2}(\tilde{\mathcal{P}},h)MD^{\prime}(\tilde{\mathcal{P}}h)K_{s(\tilde{\mathcal{P}}h)},$
@@
$D(\tilde{\mathcal{P}}h)K_{s(\tilde{\mathcal{P}}h)}=MD^{\prime}(\tilde{\mathcal{P}}h)K_{s(\tilde{\mathcal{P}}h)}.$ (40)
@@
$=$ $\omega_{2}(h,\tilde{\mathcal{P}}^{-1})D^{\prime}(h\tilde{\mathcal{P}}^{-1})K_{s(\tilde{\mathcal{P}}h)}M^{-1},$
@@
$D(h\tilde{\mathcal{P}}^{-1})K_{s(\tilde{\mathcal{P}}h)}=D^{\prime}(h\tilde{\mathcal{P}}^{-1})K_{s(\tilde{\mathcal{P}}h)}M^{-1}.$ (41)
@@
$=$ $MD^{\prime}(\tilde{\mathcal{P}}h_{1})K_{s(\tilde{\mathcal{P}}h_{1})}D^{\prime}(h_{2}\tilde{\mathcal{P}}^{-1})K_{s(\tilde{\mathcal{P}}h_{2})}M^{-1}$
$=$ $\omega_{2}(\tilde{\mathcal{P}}h_{1},h_{2}\tilde{\mathcal{P}}^{-1})MD^{\prime}(\tilde{\mathcal{P}}h_{1}h_{2}\tilde{\mathcal{P}}^{-1})K_{h_{1}h_{2}}M^{-1}.$
Denoting $\tilde{\mathcal{P}}h_{1}h_{2}\tilde{\mathcal{P}}^{-1}=g\in G_{f}(k)$, then we have $D^{\prime}(g)K_{s(g)}=MD^{\prime}(g)K_{s(g)}M^{-1}$, namely
$D(g)K_{s(g)}=MD(g)K_{s(g)}M^{-1}.$
Since $h_{1},h_{2}$ are arbitrary elements in $G_{f}(k)$, we conclude that for any $g\in G_{f}(k)$ the unitary matrix $M$ commutes with the restricted rep $D(G_{f}(k))$.
From the above lemma1, if two sets of irreducible zero modes only differ by the reps of $\tilde{\mathcal{P}}$, then the difference $M=D(\tilde{\mathcal{P}})D^{\prime-1}(\tilde{\mathcal{P}})$ must belong to the centralizer of $D(G_{f}(k))$. This is a necessary condition under which two sets of irreducible zero modes can couple with each other to form a set of reducible nonzero modes. In the following, we will list all possible cases that may lead to a set of reducible minimal nonzero modes.

### v3

@@ Appendix B Proof of $[D(\mathcal{P})K,D(g)K_{s(g)}]=0$
Now we verify equation (33). Suppose $g$ acts on $C_{\uparrow}^{\dagger}$ as
@@ Appendix C Generalized band representation for SCs/SFs/QSLs
When the pairing terms are switched off, the SCs/SFs become metals or insulators where a number of localized orbitals in the unit cell in real space determine the energy band structure in the Brillouin zone. The energy bands as entire entities form a special representation of the system’s symmetry group $G$ called the band representation Zak [1982], Bradlyn et al. [2017], Cano and Bradlyn [2021]. Such a band rep is called elementary if it cannot be decomposed as a direct sum of smaller band representations. Once the elementary band reps are known, all the band reps can be obtained by the direct sum of elementary ones.
The band representations can be generalized to BdG systems by introducing ‘particle-hole’ degrees of freedom. We start from the full ‘symmetry’ group $F_{f}=G_{f}+\mathcal{P}G_{f}$. For simplicity, here we focus on the case where the translation symmetries of $G_{f}$ are not fractionalized (i.e. the unit cell is not enlarged). The construction of an elementary band representation needs two ingredients. The first is a ‘maximal’ Wyckoff position $q_{1}$ with a site-symmetry group $G_{q_{1}}$ (composed of operations in $G_{f}$ which stabilize the site $q_{1}$). The second is a set of creation and annihilation operators of localized orbitals $\phi_{1}(q-q_{1}-R_{m}),\phi_{2}(q-q_{1}-R_{m}),...,\phi_{r}(q-q_{1}-R_{m})$ centered at $q_{1}$ in the $m$-th unit cell, namely $[c_{\sigma,q_{1}+R_{m}}^{\dagger},c_{\sigma,q_{1}+R_{m}}]$ with $\sigma=1,2,...,r$ and
@@
Moreover, if one defines the representation matrix for the little co-group element $g_{\alpha}=(s_{\alpha},c_{\alpha}||p_{\alpha})\in G_{f}(k)$ with
$D^{(l)k}_{(\Sigma,q_{i}),(\Sigma^{\prime},q_{j})}(g_{\alpha})$ $=$ $D^{(l)k}_{(\Sigma,q_{i}),(\Sigma^{\prime},q_{j})}(\alpha)e^{\mathrm{i}k\cdot t_{\alpha}}$
@@
then $D^{(l)k}_{(\Sigma,q_{i}),(\Sigma^{\prime},q_{j})}(G_{f}(k))$ form a projective rep of the little co-group $G_{f}(k)$ with the factor system
$\omega_{2}(g_{\alpha},g_{\beta})=e^{-iK_{\alpha}\cdot t_{\beta}}$
where $g_{\alpha}=(s_{\alpha},c_{\alpha}||p_{\alpha}),g_{\beta}=(s_{\beta},c_{\beta}||p_{\beta})$, $K_{\alpha}=p_{\alpha}^{-1}\cdot k-k$ with $\alpha,\beta\in G_{f}$ and $g_{\alpha},g_{\beta}\in G_{f}(k)$.
@@ Appendix D The restriction from irreps of $F_{f}(k)$ to its normal subgroup $G_{f}(k)$
We start from a theorem:
@@
If a group $G$ has normal subgroup $H$ with $G=H+rH,r^{2}\in H$ and $G/H\cong Z_{2}$, supposing a Hilbert space $\mathcal{L}$ carries a projective irrep $D(H)$ of $H$, then the restricted rep $D(H)$ contains at most two irreps of $H$. There are two possibilities:
  • (I)
    $D(H)$ is an irrep of $H$;
  • (II)
    $D(H)$ can be transformed into a direct sum of two irreps, namely $\mathcal{L}=\mathcal{L}_{1}\oplus\mathcal{L}_{2}$ where $\mathcal{L}_{1},\mathcal{L}_{2}$ each carries an irrep of $H$, and $r$ permutes $\mathcal{L}_{1},\mathcal{L}_{2}$.
@@
Since $\mathcal{L}$ carries a representation of $H$, it must contain an invariant subspace $\mathcal{L}_{1}\subseteq\mathcal{L}$ of $H$. Namely, for any $u\in H$, one has
$u\mathcal{L}_{1}=\mathcal{L}_{1}.$
@@
Meanwhile, $\mathcal{L}$ carries an irrep of $G$, so $r$ transforms $\mathcal{L}_{1}$ into another linear subspace $\mathcal{L}_{2}\equiv r(\mathcal{L}_{1})\subseteq\mathcal{L}$, then it is easily seen that $\mathcal{L}_{2}$ is also invariant under the action of $u\in H$:
$u\mathcal{L}_{2}=ur\mathcal{L}_{1}$ $=$ $r\left(r^{-1}ur\right)\mathcal{L}_{1}$
$=$ $ru^{\prime}\mathcal{L}_{1}=r\mathcal{L}_{1}=\mathcal{L}_{2},$
@@
where we have used the fact that $u^{\prime}=(r^{-1}ur)\in H$. Therefore, $\mathcal{L}_{2}$ also carries an irrep of $H$.
@@
Since both $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are irreducible representation spaces of $H$, there are two possibilities:
(I) $\mathcal{L}_{1}=\mathcal{L}_{2}=\mathcal{L}$. In this case, the restricted rep $D(H)$ from $D(G)$ is irreducible, we call such $D(G)$ a simple irrep;
(II) $\mathcal{L}_{1}\cap\mathcal{L}_{2}=\emptyset$. In this case, $\mathcal{L}^{\prime}=\mathcal{L}_{1}\cup\mathcal{L}_{2}$ is an irrep space of $G$. Noticing that $\mathcal{L}$ is an irrep space of $G$, and that irreducible space of $G$ containing $\mathcal{L}_{1}$ as a subspace is unique, we have $\mathcal{L}^{\prime}=\mathcal{L}$, namely $\mathcal{L}=\mathcal{L}_{1}\cup\mathcal{L}_{2}$. In this case, the restricted rep $D(H)$ from $D(G)$ is reducible, we call such $D(G)$ a composite irrep.
The above proof is valid no matter if $H$ is an unitary group or an anti-unitary, and no matter if the element $r$ is unitary or anti-unitary.
The structure of the simple irreps in case (I) is relatively simple since $D(H)$ is already irreducible. Here we discuss the structure of the composite irreps in case (II).
Fixing a set of bases $[\Phi]=[\phi_{1},...,\phi_{n}]$ in $\mathcal{L}_{1}$ (assuming the dimensionality of $\mathcal{L}_{1}$ is $n$), and denote $d(H)$ as the $d$-dimensional representation of $H$ on $\mathcal{L}_{1}$. In the other subspace $\mathcal{L}_{2}$, we simply choose the bases as $r[\Phi]=[r\phi_{1},...,r\phi_{n}]$. Then in the space $\mathcal{L}$ the particle-hole operator $r$ is represented as
$D(r)K_{s(r)}=\left(\begin{matrix}0&\omega_{2}(r,r)d(r^{2})\\ I_{n}&0\end{matrix}\right)K_{s(r)},$
where $r^{2}\in H$. For any $u\in H$, we have
$D(u)K_{s(u)}=\begin{pmatrix}d(u)&0\\ 0&d^{\prime}(u)\end{pmatrix}K_{s(u)}.$
with
$d^{\prime}(u)={\omega_{2}(u,r)\over\omega_{2}(r,r^{-1}gr)}K_{s(r)}d(r^{-1}ur)K_{s(r)}.$
Now we apply the above theorem to the little co-group $F_{f}(k)$ at momentum $\bm{k}$ that is invariant under $\tilde{\mathcal{P}}$, with $F_{f}(k)=G_{f}(k)+\tilde{\mathcal{P}}G_{f}(k)$ and
$F_{f}(k)/G_{f}(k)\cong Z_{2}.$
Denoting $D(F_{f}(k))$ as a projective irrep of $F_{f}(k)$, according to theorem 5 the restricted rep $D(G_{f}(k))$ is either irreducible or a direct sum of two irreps. In the following we discuss three different situations of $F_{f}(k)$.
(1) $F_{f}(k)$ is an unitary group. The simple irrep of $F_{f}(k)$ either contribute a set of irreducible zero modes (if there are effective charge conjugation symmetry) or couple with another simple irrep to form a set of RMNZM; the composite irrep of $F_{f}(k)$ gives rise to a set of irreducible non-zero modes, where the energy eigenspaces $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ have opposite energies ±$\pm\varepsilon_{k}$.
@@
(2) $F_{f}(k)$ is an anti-unitary group, $G_{f}(k)$ is unitary and $\tilde{\mathcal{P}}$ is anti-unitary. The irrep of $F_{f}(k)$ falls into three classes, namely the $\mathbb{R},\mathbb{C},\mathbb{H}$ classes. The irrep in the $\mathbb{R}$ class is a simple irrep; the $\mathbb{C}$ and $\mathbb{H}$ classes are composite irreps (irreducible nonzero modes), the two classes are distinguished by whether the representations carried by $\mathcal{L}_{1}$ is equivalent to that of $\mathcal{L}_{2}$ or not)Shaw and Lever [1974].
@@
(3) $F_{f}(k)$ is anti-unitary, $G_{f}(k)$ is anti-unitary and $\tilde{\mathcal{P}}$ is unitary. Similar to the case where $F_{f}(k)$ is unitary, the irrep of $F_{f}(k)$ can be simple or composite.
@@ Appendix E Reducible minimal nonzero modes
A set of reducible minimal nonzero modes (RMNZM) contains two simple irreps of $F_{f}(k)$. To ensure that the two irreps can couple with each other to form nonzero modes, the restricted reps of the normal subgroup $G_{f}(k)$ should be equivalent. Now we start with two irreps $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ satisfying the relation
@@
Defining a matrix
$X\equiv D(\tilde{\mathcal{P}})D^{\prime-1}(\tilde{\mathcal{P}}),$
then we have the following lemma:
@@
If two simple irreps $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ satisfies the relation $D(G_{f}(k))=D^{\prime}(G_{f}(k))$, then the unitary matrix $X=D(\tilde{\mathcal{P}})D^{\prime-1}(\tilde{\mathcal{P}})$ commutes with the restricted rep $D(G_{f}(k))$.
@@
$=$ $XD^{\prime}(\tilde{\mathcal{P}})D^{\prime}(h)K_{s(\tilde{\mathcal{P}}h)}$
$=$ $\omega_{2}(\tilde{\mathcal{P}},h)XD^{\prime}(\tilde{\mathcal{P}}h)K_{s(\tilde{\mathcal{P}}h)},$
@@
$D(\tilde{\mathcal{P}}h)K_{s(\tilde{\mathcal{P}}h)}=XD^{\prime}(\tilde{\mathcal{P}}h)K_{s(\tilde{\mathcal{P}}h)}.$ (40)
@@
$=$ $\omega_{2}(h,\tilde{\mathcal{P}}^{-1})D^{\prime}(h\tilde{\mathcal{P}}^{-1})K_{s(\tilde{\mathcal{P}}h)}X^{-1},$
@@
$D(h\tilde{\mathcal{P}}^{-1})K_{s(\tilde{\mathcal{P}}h)}=D^{\prime}(h\tilde{\mathcal{P}}^{-1})K_{s(\tilde{\mathcal{P}}h)}X^{-1}.$ (41)
@@
$=$ $XD^{\prime}(\tilde{\mathcal{P}}h_{1})K_{s(\tilde{\mathcal{P}}h_{1})}D^{\prime}(h_{2}\tilde{\mathcal{P}}^{-1})K_{s(\tilde{\mathcal{P}}h_{2})}X^{-1}$
$=$ $\omega_{2}(\tilde{\mathcal{P}}h_{1},h_{2}\tilde{\mathcal{P}}^{-1})XD^{\prime}(\tilde{\mathcal{P}}h_{1}h_{2}\tilde{\mathcal{P}}^{-1})K_{h_{1}h_{2}}X^{-1}.$
Denoting $\tilde{\mathcal{P}}h_{1}h_{2}\tilde{\mathcal{P}}^{-1}=g\in G_{f}(k)$, then we have $D^{\prime}(g)K_{s(g)}=XD^{\prime}(g)K_{s(g)}X^{-1}$, namely
$D(g)K_{s(g)}=XD(g)K_{s(g)}X^{-1}.$
Since $h_{1},h_{2}$ are arbitrary elements in $G_{f}(k)$, we conclude that for any $g\in G_{f}(k)$ the unitary matrix $X$ commutes with the restricted rep $D(G_{f}(k))$.
From the above lemma1, if two simple irreps of $F_{f}(k)$ only differ by the reps of $\tilde{\mathcal{P}}$, then the difference $X=D(\tilde{\mathcal{P}})D^{\prime-1}(\tilde{\mathcal{P}})$ must belong to the centralizer of $D(G_{f}(k))$. This is a necessary condition under which two simple irreps can couple with each other to yield a set of reducible nonzero modes. In the following, we will list all possible cases that may lead to a set of RMNZM.